40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 12     Entry time: Wed Oct 24 08:58:09 2007
Author: steve 
Type: Other 
Category: PSL 
Subject: laser headtemp is up 
C1:PSL-126MOPA_HTEMP is 19.3C

Half of the chiller's air intake was covered by loose paper
Attachment 1: htempup.jpg  90 kB  | Hide | Hide all
htempup.jpg
ELOG V3.1.3-